The following simple random sample was selected from a normal distribution: 4, 6, 3, 5, 9, and 3.

a.     Construct a 90% confidence interval for the population mean μ.

 

b.    Construct a 95% confidence interval for the population mean μ.

 

c.     Construct a 99% confidence interval for the population mean μ.


1
Expert's answer
2022-04-20T00:11:30-0400
"mean=\bar{x}=\dfrac{1}{6}(4+6+3+5+9+3)=5""s^2=\dfrac{1}{6-1}((4-5)^2+(6-5)^2+(3-5)^2+(5-5)^2""+(9-5)^2+(3-5)^2)=5.2"





"s=\sqrt{s^2}=\sqrt{5.2}\approx2.280351"

1. The critical value for "\alpha = 0.1" and "df = n-1 = 5"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.015036." The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(5-2.015036\times\dfrac{2.280351}{\sqrt{6}},""5+2.015036\times\dfrac{2.280351}{\sqrt{6}})""=(3.1241, 6.8759)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "3.1241 < \mu < 6.8759," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(3.1241, 6.8759)."


2. The critical value for "\alpha = 0.05" and "df = n-1 = 5"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.570543." The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(5-2.570543\times\dfrac{2.280351}{\sqrt{6}},""5+2.570543\times\dfrac{2.280351}{\sqrt{6}})""=(2.6070, 7.3930)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "2.6070 < \mu < 7.3930," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(2.6070, 7.3930)."


3. The critical value for "\alpha = 0.01" and "df = n-1 = 5"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =4.031677." The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(5-4.031677\times\dfrac{2.280351}{\sqrt{6}},""5+4.031677\times\dfrac{2.280351}{\sqrt{6}})""=(1.2467, 8.7533)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "1.2467 < \mu < 8.7533," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(1.2467, 8.7533)."


4.

a. The critical value for "\alpha = 0.1" and "df = n-1 = 24"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =1.710882." The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(5-1.710882\times\dfrac{2.280351}{\sqrt{25}},""5+1.710882\times\dfrac{2.280351}{\sqrt{25}})""=(4.2197, 5.7803)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "4.2197 < \mu < 5.7803," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(4.2197, 5.7803)."


b. The critical value for "\alpha = 0.05" and "df = n-1 = 24"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.063899." The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(5-2.063899\times\dfrac{2.280351}{\sqrt{25}},""5+2.063899\times\dfrac{2.280351}{\sqrt{25}})""=(4.0587, 5.9413)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "4.0587 < \mu < 5.9413," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(4.0587, 5.9413)."


c. The critical value for "\alpha = 0.01" and "df = n-1 = 24"  degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.79694." The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(5-2.79694\times\dfrac{2.280351}{\sqrt{25}},""5+2.79694\times\dfrac{2.280351}{\sqrt{25}})""=(3.7244, 6.2756)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "3.7244 < \mu < 6.2756," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(3.7244, 6.2756)."


The width of a confidence interval decreases as the sample size increases and increases as the confidence level increases.

Larger samples give narrower intervals. We are able to estimate a population mean more precisely with a larger sample size.